# prompt[Prompt][List of Prompts]

### v1

@@ Abstract
Large language models (LLMs) face significant challenges with needle-in-a-haystack tasks, where relevant information (“the needle“) must be drawn from a large pool of irrelevant context (“the haystack“). Previous studies have highlighted positional bias and distractor quantity as critical factors affecting model performance, yet the influence of gold context size has received little attention. We address this gap by systematically studying how variations in gold context length impact LLM performance on long-context question answering tasks. Our experiments reveal that LLM performance drops sharply when the gold context is shorter, i.e., smaller gold contexts consistently degrade model performance and amplify positional sensitivity, posing a major challenge for agentic systems that must integrate scattered, fine-grained information of varying lengths. This pattern holds across three diverse domains (general knowledge, biomedical reasoning, and mathematical reasoning) and seven state-of-the-art LLMs of various sizes and architectures. Our work provides clear insights to guide the design of robust, context-aware LLM-driven systems.
Figure 1: Changing both the size and position of gold context (relevant information) within a set of distracting context(irrelevant information), we observe that LLM Needle-In-A-Haystack Performance is overall lower and more sensitive to position when models are given short gold documents (dashed line) as opposed to long (solid line).
@@ 1 Introduction
Large language models (LLMs) increasingly power applications that require reasoning over vast amounts of information, ranging from synthesizing findings across scientific literature  [1, 2, 3, 4, 5, 6, 7], to navigating complex codebases  [8, 9, 10], to maintaining coherence in multi-turn conversations. These applications share a common requirement: strong long-context understanding. This is particularly vital for agentic systems, in which autonomous agents must integrate heterogeneous information streams from specialized components to reason, plan, and act effectively.
@@
A critical stage in such systems is aggregation, the synthesis of retrieved evidence into an accurate, actionable response. This stage determines what content to include, cite, or ignore, and has direct implications for safety, reliability, and factual correctness. Aggregation becomes especially challenging in needle-in-a-haystack scenarios, where relevant evidence (the ‘gold context’) is embedded within a large volume of topically related or superficially plausible but ultimately irrelevant or misleading, ‘distractor context’  [11, 12]. Successful aggregation requires precise identification and prioritization of minimal but essential content while discarding noisy signals.
Although LLMs now support context windows stretching into the millions of tokens, recent studies show that simply increasing input length does not ensure strong long-context reasoning. Prior work has explored positional bias  [13, 14, 15], showing that early content is more likely to be attended to, and that distractors degrade performance. However, one key dimension remains underexplored: how the size of the gold context influences model performance?
In this study, we systematically examine how gold context size affects long-context performance in LLMs. Specifically, we adapt three diverse benchmarks, CARDBiomedBench  [16] (biomedical reasoning), NaturalQuestions  [17] (general knowledge), and NuminaMath1.5  [18] (mathematical reasoning), to vary both the size and position of the gold context while keeping the distractor content fixed (see Section 2).
Our key findings are:
  1. 1.
    Smaller gold contexts lead to significantly worse performance.
  2. 2.
    Smaller gold contexts exhibit higher positional sensitivity.
Models are remarkably more sensitive to the placement of smaller gold spans, with accuracy declining when relevant content appears later in the context window. This effect holds across seven state-of-the-art open and closed-weight LLMs spanning diverse architectures and scales. Notably, models achieve near-perfect scores in no-distractor settings, confirming that these failures are due to aggregation breakdowns rather than task difficulty (see Section 3).
Further exploration provides additional insights into the nuances of aggregation (Section 4). We observe a pronounced primacy bias for smaller gold contexts, wherein relevant information placed near the beginning of the context window is more reliably utilized by models. This effect diminishes with larger gold contexts, which are more resilient to positional variation. The problem is amplified in domain-specific reasoning (e.g., biomedical and mathematical tasks), and exacerbated as distractor size increases. We also find that domain-specific tasks exacerbate the challenges posed by small gold contexts relative to general-knowledge tasks. Last but not least, we observe that this trend persists as number of distractor documents increases.
These findings carry practical implications. In real-world deployments, context size, position, distractor size, and noise are rarely controllable. Aggregation failures due to overlooked gold context size can degrade trust, safety, and downstream task performance. Based on our empirical findings, we offer guidelines for designing more robust aggregation strategies for agentic systems, including techniques for structuring and expanding critical evidence to reduce fragility.
In sum, we show that the size of relevant evidence—not just its location—is a critical factor in long-context reasoning. Our results highlight an often-overlooked bottleneck in LLM capabilities and offer both diagnostic insights and prescriptive tools for practitioners building context-resilient AI systems.
@@ 2 Experimental Setup
We designed our experiments to systematically evaluate how the size of the gold context affects long-context LLM performance. This section outlines our core design objectives, benchmark adaptations, baseline validations, and primary evaluations simulating realistic aggregation settings.
@@ 2.1 Desiderata
To systematically evaluate the impact of gold context size on long-context LLM aggregation performance, our guiding desiderata were (1) Realism, (2) Gold Size Variability, (3) Distractors, and (4) Generality.
Realism.
In real-world agentic systems, aggregation typically involves synthesizing outputs from multiple specialized agents, each retrieving documents from their domain of expertise. Usually, one agent returns the document containing the correct answer (the “gold” document), while others provide distractors, topically relevant but ultimately uninformative documents. We simulate this by inserting a gold document of varying size at different positions within a fixed-length sequence of distractors. The document order is randomized to reflect natural uncertainty in agent contributions and retrieval quality.
Gold Size Variability.
To isolate the impact of gold context size, we constructed three nested variants for each benchmark:
@@
These were wrapped as pseudo-documents, including metadata such as titles and questions when appropriate. All variants were hierarchically structured (small $\subset$ medium $\subset$ large) and validated for sufficiency. See  2 for examples. Performance is very high and uniform when observing solely any size of gold as can be seen in Table 1.
Distractors.
To simulate realistic confounding signals, we curated distractors that are topically relevant and lexically similar to the question but do not contain the answer. Distractor quantities per benchmark were calibrated to match token distributions observed in a real-world multi-agent retrieval system ($\sim 20k$ tokens).
Generality.
We selected three diverse benchmarks spanning biomedical, general knowledge, and mathematical reasoning, and evaluated performance across seven leading LLMs of varying architecture and scale. This ensures that our findings generalize across domains and model classes.
@@ 2.2 Task Construction: Needles and Haystacks
We adapted three established question and answering benchmarks—CARDBiomedBench (biomedical reasoning), NaturalQuestions (general knowledge), and NuminaMath1.5 (mathematical reasoning)—to create controlled needle-in-a-haystack settings. Gold context sizes were varied, accompanied by distractors explicitly designed to be topically relevant yet answer-free. Figure 3 displays token count distributions for the varying sizes of gold.
@@
Figure 3: Token count distributions for the varying sizes of gold context on each benchmark. Median token count is in parenthesis in the legend. X-axis is dynamically scaled as linear (0-500 tokens) and logarithmic (500+ tokens) for visibility.
CARDBiomedBench (CBB).
CBB is a question-answering dataset focused on neurodegenerative diseases (NDDs), designed to evaluate LLM performance on biomedical reasoning tasks involving genetic, molecular, and clinical information. The BiomedSQL  [19] variant augments each example with SQL queries and database rows to support structured reasoning experiments. This provides the following sizes of gold:
Small: Answer only. Medium: SQL query + answer. Large: SQL query + returned rows + answer.
Distractors were sampled from documents retrieved by four independent domain-specific agents in a real-world system. These documents are semantically relevant but verifiably do not contain the answer, presenting realistic aggregation challenges.
NaturalQuestions (NQ).
NQ is an open-domain QA benchmark using real Google user queries, with evidence-linked Wikipedia passages from the KILT corpus  [20]. Gold contexts were derived as:
Small: Sentence containing the answer. Medium: Full paragraph around the sentence. Large: Paragraph $\pm$ 4 adjacent paragraphs.
Distractors were drawn from the HELMET  [21] adaptation of NQ-KILT, using dense retrieval over 2019 Wikipedia with the gte-large-en-v1.5 embedding model  [22]. From HELMET’s 100-token chunks, we excluded documents: (1) explicitly labeled as gold evidence and (2) containing the answer string to ensure high semantic similarity without answer leakage. This ensured the distractors are lexically and topically aligned with the question, but verifiably devoid of answer content.
NuminaMath1.5 (NM).
NM is the largest open-source dataset for math reasoning, with problems ranging from high school to International Mathematical Olympiad (IMO)-level difficulty, originating from diverse sources like Chinese K-12 exams, AMC/AIME contests, and global math forums. We used the OpenR1Math  [23] variant, which includes model-generated solution traces from DeepSeekR1  [24] verified for correctness. We filtered for examples with complete reasoning streams and defined gold variants as:
Small: Final answer. Medium: Textbook-style solution + answer. Large: Full LLM-generated chain-of-thought + solution + answer.
Distractors were unrelated reasoning traces matched for length and complexity. Due to length variability, large gold contexts were capped at the final 5k tokens, which included concluding reasoning and answers.
@@ 2.3 Baseline Experiments
To validate our construction, we ran three baseline conditions:
Closed-book.
No context was provided, assessing whether models could answer from internal knowledge. This gauges possible benchmark saturation.
Gold-only.
Each gold context size (small, medium, large) was presented alone, without distractors. This verified that all variants were independently sufficient to solve the task and that downstream performance drops are due to aggregation effects (e.g., distractor interference or gold placement). Table  1 contains a subset of such baselines.
Distractor-only.
Models were given only distractor documents. For CBB, we also tested distractors from each agent separately to confirm they were individually non-informative. These checks ensured that distractors lacked sufficient signal to answer correctly.
These baselines serve to validate that observed performance differences in main experiments result from changes to gold size, rather than underlying flaws in dataset or distractor construction.
CBB NQ NM
small large small large small large
GPT-4o 0.98 1.00 0.70 0.74 0.93 0.88
Gemini-2.0-Flash 0.96 0.96 0.86 0.80 0.97 0.94
LLaMA-3.1-405b 0.98 1.00 0.81 0.80 0.90 0.92
LLaMA-3.3-70b 0.98 1.00 0.82 0.80 0.89 0.92
Table 1: Baseline accuracy when models are presented with only the gold context (small or large). Results confirm that all gold variants are independently sufficient to solve the task. Minor fluctuations reflect benchmark variability rather than consistent advantages for larger gold contexts.
@@ 2.4 Main Experiments
We simulated realistic aggregation scenarios by embedding each gold context size at varying positions within a fixed sequence of distractors. This tests both gold size sensitivity and positional sensitivity simultaneously.
We evaluated seven leading LLMs:
  • Closed-weight: GPT-4o  [25], GPT-4o-Mini  [26], Gemini-2.0-Flash, Gemini-2.0-Flash-Lite  [27]
  • Open-weight: LLaMA-3.1-405B, LLaMA-3.3-70B, LLaMA-3.1-8B  [28]
Each model was evaluated on every size-position combination in a controlled, deterministic setting (temperature = 0). Prompts were standardized within each benchmark. This setup enables rigorous, cross-model evaluation of gold context sensitivity and simulates the type of unpredictable document ordering common in deployed LLM systems.
3 Empirical Findings
Our experiments reveal that the size of the gold context has a substantial and consistent effect on long-context performance across benchmarks, models, and domains.
3.1 Smaller Gold Contexts Lead to Lower Performance
Across all benchmarks and models, increasing the size of the gold context significantly improves accuracy (Figure 4). For example, on CARDBiomedBench (CBB), Gemini-2.0-Flash improved from 48% with small gold to 62% with medium and 73% with large gold contexts. GPT-4o showed a similar trend, rising from 77% (small) to 98% (large), while LLaMA-3.1-405B improved from 74% to 92%.
Notably, performance with large gold contexts closely approached the model-specific gold-only baselines (i.e., accuracy when the gold context is shown without any distractors) recorded at 96% for Gemini-2.0-Flash, and 100% for both GPT-4o and LLaMA-3.1-405B. This suggests that large gold contexts allow models to nearly recover ideal aggregation performance, while small golds fall significantly short.
@@ 3.1 Smaller Gold Contexts Lead to Lower Performance
3.2 Smaller Gold Contexts Are More Sensitive to Position
We observed a clear positional sensitivity effect that was amplified in smaller gold contexts (Figure 5). Performance systematically declined when small gold contexts appeared later in the input, while large gold contexts were more robust to position.
For instance, in CBB, Gemini-2.0-Flash achieved 94% accuracy when the small gold context was placed at the start of the context window, but only 33% when placed near the end, a 61-point drop. In contrast, the large gold context declined more gradually, from 84% to 65%, demonstrating greater positional resilience. This pattern held across all evaluated models and benchmarks.
Importantly, the positional effect was more pronounced in domain-specific tasks (e.g., biomedical and math) than in general knowledge (NQ), suggesting that both information type and gold size compound aggregation difficulty.
@@ 3.2 Smaller Gold Contexts Are More Sensitive to Position
These empirical results establish that both gold context size and position critically affect long-context reasoning. In the next section, we delve deeper into these patterns, analyzing where, why, and how these failures manifest at finer granularity.
4 Additional Analysis
Beyond our core findings, we conduct further analyses to understand why smaller gold contexts lead to performance degradation, and under which conditions these effects are most severe. Specifically, we examine the impact of domain specificity, positional variance, and distractor volume.
Domain-Specific Tasks Amplify Sensitivity to Gold Context Size. The effects of gold context size are notably amplified in domain-specific tasks compared to general knowledge. Figure 6 quantifies this by measuring the range in model performance across different gold context positions. For each model and gold size, we compute the performance range as the difference between maximum and minimum scores across all positions:
$\text{Range}=\mathop{\max}_{i\in\{1,\dots,n\}}\text{perf}(\text{position}_{i})-\mathop{\min}_{i\in\{1,\dots,n\}}\text{perf}(\text{position}_{i})$ (1)
(a) CARDBiomedBench
(b) NaturalQuestions
(c) NuminaMath1.5
Figure 6: Positional sensitivity by benchmark. For each model and gold context size, we compute the range (Equation  1) of performance across positions. Smaller gold contexts exhibit much higher sensitivity (larger ranges), especially in domain-specific tasks (CBB, NM). Larger gold contexts yield more stable performance across positions.
Smaller Gold Contexts Exhibit Stronger Primacy Bias. We also observed a primacy bias across models, performance was consistently higher when the gold context appeared early in the input window. This effect was especially pronounced for small gold contexts. In some cases, small gold contexts placed at the beginning of the input even outperformed medium or large contexts placed later, despite their reduced information content. This occurs often in the left and right columns of Figure 5, where the small gold line starts at the top at gold position 0.0 before crossing over to the bottom.

### v2

@@ Abstract
Large language models (LLMs) face significant challenges with needle-in-a-haystack tasks, where relevant information (“the needle“) must be drawn from a large pool of irrelevant context (“the haystack“). Previous studies have highlighted positional bias and distractor quantity as critical factors affecting model performance, yet the influence of gold context size, the length of the answer-containing document, has received little attention. We present the first systematic study of gold context size in long-context question answering, spanning three diverse benchmarks (general knowledge, biomedical reasoning, and mathematical reasoning), eleven state-of-the-art LLMs (including recent reasoning models), and more than 150K controlled runs. Our experiments reveal that LLM performance drops sharply when the gold context is shorter, i.e., smaller gold contexts consistently degrade model performance and amplify positional sensitivity, posing a major challenge for agentic systems that must integrate scattered, fine-grained information of varying lengths. This effect persists under rigorous confounder analysis: even after controlling for gold context position, answer token repetition, gold-to-distractor ratio, distractor volume, and domain specificity, gold context size remains a decisive, independent predictor of success. Our work provides clear insights to guide the design of robust, context-aware LLM-driven systems.
@@ 1 Introduction
Large language models (LLMs) increasingly power applications that require reasoning over vast amounts of information, such as synthesizing evidence across scientific literature (sprueill2024chemreasoner; bazgir2025agentichypothesis; weiqiwang2025arxiv2table; gao2025sciencehierarchy) or navigating complex codebases (liu2023repobench; zhang2023repocoder; bogomolov2024long).
@@
A critical stage in such systems is aggregation, the synthesis of retrieved evidence into an accurate, actionable response. This stage determines what content to include, cite, or ignore, and has direct implications for safety, reliability, and factual correctness. One specific, common variant of aggregation is needle-in-a-haystack (NIAH) scenarios, where relevant evidence (the ‘gold context’) is embedded within a large volume of topically related or superficially plausible but ultimately irrelevant or misleading, ‘distractor context’ (tay2021long; shaham2022scrolls). Successful aggregation requires precise identification of essential evidence among a large number of distractor documents.
In NIAH scenarios, prior work has explored phenomena such as positional bias (wang2023primacy; liu2024lost), showing that models disproportionately attended to early content and that distractors degrade performance. Yet one key dimension remains underexplored: how the size of the gold context influences model performance. This question is critical for the design of agentic systems, where autonomous agents must perform NIAH-style aggregation over heterogeneous information streams produced by specialized components, whose evidence can vary widely in length.
In this study, we present a systematic analysis of gold context size as an independent variable in LLM performance in long-context, single-needle NIAH settings. Our main finding (§3) is that smaller gold contexts consistently (1) degrade performance and (2) heighten sensitivity to positional bias, as illustrated in Fig.1. This exposes a form of brittleness not captured in prior work.
Figure 1: LLM performance on a needle-in-a-haystack task (CARDBiomedBench). The y-axis is a measure of accuracy (higher is better). Varying both the size and position of the gold context within fixed distractor content, we find that LLMs perform worse and exhibit greater positional sensitivity when given short gold contexts (dashed line) compared to long ones (solid line).
To support our analysis, we adapt three diverse datasets spanning mathematics, biomedical reasoning, general knowledge, systematically varying both the size and position of the gold context while keeping the distractor content fixed2). Our study comprises over 150K controlled runs across eleven state-of-the-art LLMs, both general-purpose and reasoning-focused, ensuring that the observed effects hold across architectures and tasks.
We conduct extensive confounder analyses to ensure that our findings are not driven by spurious factors (§4). Even after controlling for gold position, answer-token repetition, gold-to-distractor ratio, distractor volume, and domain specificity, gold context size remains a strong, independent predictor of performance. The overall pattern is consistent—larger gold contexts outperform smaller ones—though the magnitude of the effect varies across conditions. For instance, smaller gold contexts are particularly vulnerable to primacy bias, whereas larger gold ones exhibit greater robustness to both positional variation and distractor interference.
These findings have practical implications. In real-world deployments, factors such as context size, position, distractor length, and noise are typically uncontrolled. Our results show that heterogeneity in document length can induce aggregation failures, particularly when small but critical evidence is overshadowed by much longer passages. This situation naturally arises in retrieval-augmented and agentic systems, where an aggregator must isolate a single relevant piece of evidence from a mixture of heterogeneous sources. Practitioners can mitigate this fragility by monitoring length disparities among retrieved documents and by avoiding pipelines that combine extremely short and very long contexts, thereby reducing size-induced attention asymmetries.
In summary, our contributions are as follows:
  • Novel determinant of long-context performance: To the best of our knowledge, we are the first to demonstrate that the size of gold contexts functions as a hidden variable in LLM performance on long-context NIAH. Smaller golds degrade accuracy and amplify positional bias, underscoring a potential fragility in real-world applications.
  • Robust to confounders: We identify and analyze five potential confounding variables, (1) gold context position, (2) answer token repetition, (3) gold-to-distractor ratio, (4) distractor volume, and (5) domain specificity, demonstrating that gold context size remains a decisive predictor of success despite these factors.
  • Large-scale experimentation: We repeat our experiments and aggregate findings across eleven state-of-the-art LLMs, three diverse benchmarks, three sizes of gold, and six positions in the context window totaling over 150k controlled runs.
@@ 2 Experimental Setup
This section outlines our design objectives, benchmark adaptations, baseline validations.
@@ 2.1 Desiderata
We begin by outlining the core desiderata guiding our experimental design.
Realism. In real-world agentic systems, aggregation involves synthesizing outputs from multiple specialized agents, each retrieving information from their domain of expertise. Usually, one agent returns the document containing the correct answer (the “gold context"), while others provide distractors, topically relevant but ultimately uninformative. We simulate this by inserting a gold context of varying size at different positions within a fixed-length sequence of distractors. Document order is randomized to reflect natural uncertainty in agent contributions and retrieval quality.
Gold Size Variability. We constructed three nested gold variants for each benchmark:
@@
These were wrapped in pseudo-documents (titles, questions). Variants are hierarchically structured (small $\subset$ medium $\subset$ large) and validated for sufficiency. See Figure 9 for examples. Performance is high and uniform when observing only the gold of any size (Appendix LABEL:apdx:baselines).
Distractors. To simulate realistic scenarios, we curate distractors topically relevant and lexically similar to the question but lacking the answer. The total distractor budget per benchmark is $\sim 20k$ tokens—a deliberate design choice that provides a sufficiently long distractor context to be considered ‘long-context’ while remaining computationally manageable for extensive experimentation.
Generality. We select three diverse benchmarks spanning biomedical, general knowledge, and mathematical reasoning, and evaluate performance across eleven leading LLMs of varying architecture and scale. This ensures that our findings generalize across domains and model classes.
@@ 2.2 Task Construction: Needles and Haystacks
We adapt three established question and answering benchmarks—CARDBiomedBench (biomedical reasoning), NaturalQuestions (general knowledge), and NuminaMath1.5 (mathematical reasoning)—to create controlled NIAH settings. Gold context sizes were varied, accompanied by distractors explicitly designed to be topically relevant yet answer-free. Figure 2 displays token count distributions for the varying sizes of gold. We provide further details on datasets and their metrics in §B.1.
@@
Figure 2: Token count distributions for varying sizes of gold context on each benchmark. Median token count is in parenthesis in the legend. X-axis is scaled as linear (0-500) and logarithmic (500+).
CARDBiomedBench (CBB). CBB is a question-answering dataset focused on neurodegenerative diseases, designed to evaluate LLM performance on biomedical reasoning tasks involving genetic, molecular, and clinical information. The BiomedSQL (biomedsql_nih_card_2025) variant augments each example with SQL queries and database rows to support structured reasoning experiments:
Full answer
$\subset$ SQL query + answer $\subset$ SQL query + rows + answer
Distractors were drawn from documents retrieved by four independent domain-specific agents with access to biomedical knowledge bases. These documents are semantically relevant but verifiably do not contain the answer, presenting realistic aggregation challenges.
NaturalQuestions (NQ). NQ is an open-domain QA benchmark of Google queries, with Wikipedia passages from the KILT corpus as evidence (petroni2021kiltbenchmarkknowledgeintensive):
Sentence with the answer
$\subset$ Paragraph with the sentence $\subset$ Paragraph $\pm$ 4 paragraphs
Distractors were drawn from the HELMET (yen2025helmetevaluatelongcontextlanguage) adaptation of NQ-KILT as 100-token segments. We exclude any documents labeled as evidence or containing the answer. This ensures distractors remain lexically and topically aligned with the question, but free of the answer.
NuminaMath1.5 (NM). NM is the largest open-source dataset for math reasoning, with problems ranging from high school to International Mathematical Olympiad (IMO)-level difficulty, originating from diverse sources like Chinese K-12 exams, AMC/AIME contests, and global math forums. We used the OpenR1Math (openr1math220k) variant, which includes model-generated solution traces from DeepSeekR1 (deepseekai2025deepseekr1incentivizingreasoningcapability) verified for correctness. We filter for examples with complete reasoning streams and define gold variants as:
Full answer
[…]
Distractors were reasoning traces to different questions. Due to length variability, we cap large gold contexts at the final $5k$ tokens, which include concluding reasoning and answers.
@@ 2.3 Baseline Experiments
We run three baseline conditions to validate that observed performance differences in main experiments result from changes to gold size, rather than underlying flaws in datasets or distractor construction. Baseline results across all benchmarks and models can be found in Appendix LABEL:apdx:baselines:
  • Closed-book. No context is provided, assessing whether models could answer from internal knowledge. This gauges possible benchmark saturation.
  • Gold-only. Each gold context is presented alone, without distractors. This confirms variants were sufficient to solve the task and that downstream performance drops are due to aggregation effects (e.g., distractor interference or gold placement).
  • Distractor-only. Models are given only distractor documents. For CBB, we also test distractors from each agent separately to confirm they were individually non-informative. These checks ensure that distractors lack sufficient signal to answer correctly.
@@ 2.4 Main Experiments
We simulate realistic aggregation scenarios by embedding each gold context size at varying positions within a fixed sequence of distractors. This tests both gold size and positional sensitivity simultaneously. We evaluate eleven leading LLMs:
  • Closed-weight: o3-mini, GPT-4o (openai2024gpt4ocard), GPT-4o-Mini (openai2024gpt4omini), Gemini-2.5-Flash, Gemini-2.0-Flash, and Gemini-2.0-Flash-Lite (gemini2025flash)
  • Open-weight: DeepSeek-R1 (deepseekai2025deepseekr1incentivizingreasoningcapability), Phi-4-reasoning (abdin2025phi4reasoningtechnicalreport), LLaMA-3.1-405B, LLaMA-3.3-70B, and LLaMA-3.1-8B (dubey2024llama3herdmodels)
We evaluate each model on every size-position combination in a deterministic setting. Prompts were standardized within each benchmark. This enables rigorous, cross-model evaluation of gold context sensitivity and simulates the type of unpredictable document ordering in LLM systems.
@@
3 Main Finding: Smaller Gold Contexts Lower the Performance
Our experiments reveal that gold context size has a substantial and consistent effect on long-context performance, irrespective of confounding variables, across different benchmarks and models.
Increasing the size of the gold context significantly improves accuracy (Figure 3). On CBB, Gemini-2.0-Flash went from 48% with small to 62% with medium and 73% with large. GPT-4o performs similarly, rising from 77% (small) to 98% (large), while LLaMA-3.1-405B went from 74% to 92%.
Notably, performance with large gold contexts approaches the gold-only baselines (i.e., accuracy when the gold context is shown without any distractors) recorded at 96% for Gemini-2.0-Flash, and 100% for both GPT-4o and LLaMA-3.1-405B. This suggests that large gold contexts allow models to nearly recover ideal aggregation performance, while small golds fall significantly short.
4 Analysis of Confounding Factors
Our goal is to isolate the effect of gold context size as an independent factor in LLM performance. However, there are various other factors that confounded or correlated with our target variable, making the attribution of the observed effect non-trivial. To better understand the underlying phenomenon here, one must control for any potential confounding factors that may impact the outcome of the findings. Specifically, in this section we study and control the following confounds: gold context position in the context window (§4.1), the total number of repetitions of the answer in the context (§4.2), relative length of gold context to the total distractor evidence length (§4.3), total distractor length (§4.4) and domain specificity (§4.5).
@@ 4 Analysis of Confounding Factors
4.1 Gold Context Position
From the results in Figure 4, we observe that smaller gold contexts are hard to find regardless of their position. Nevertheless, certain positions amplify the bias against smaller gold contexts. Performance systematically declines when small gold contexts appear later in the input, while large gold contexts are more robust to position (Full results in Appendix C.1).
For instance, in CBB, Gemini-2.0-Flash achieves 94% accuracy when the small gold context is placed at the start of the context window, but only 33% when placed near the end, a 61-point drop. In contrast, the large gold context declines more gradually, from 84% to 65%, demonstrating greater positional resilience. This pattern held across all evaluated models and benchmarks with some effects more amplified than others.
Importantly, the positional effect is more pronounced in domain-specific tasks (CBB and NM) than in general knowledge (NQ), suggesting that task type and gold size compound aggregation difficulty.
We also observe that models exhibit stronger primacy bias with smaller gold contexts: performance is consistently higher when the gold context appears early in the input window. This effect is especially pronounced for small gold contexts. In some cases, small gold contexts placed at the beginning of the input even outperformed medium or large contexts placed later. This occurs often in the left and right columns of Figure 4, where the small gold line starts at the top at gold position 0.0 before crossing to the bottom.